Activity takeBatakeAssignment do?iocator sessionLo assignment ent takeCowalen Use the References to access important values if eeded for this question. How many milliliters of7.14 M perchloric acid solution should be used to prepare 3.00 L of0.300 M HCIO? 9 more group attempts remaining Retry Entire Group
Answer #1

use dilution formula

M1*V1 = M2*V2

1---> is for stock solution

2---> is for diluted solution

Use:

M1*V1 = M2*V2

V1 = (M2 * V2) / M1

V1 =(0.300*3.000)/7.14

V1 =0.126 L

   = 126 mL

Answer: 126 mL